OpenStudy (anonymous):

7-y-y= -1 what does y equal?

OpenStudy (zrsmith):

7-2y = -1 (-y-y is the same as (-y) + (-y) = -2y) 7-2y-7 = -1-7 (subtract -7 from both sides) -2y = -8 -2y/-2 = -8/-2 (divide both sides by -2) y = 4